Quentin was challenged to a half-court shooting completion. For every half-court shot he makes, he will earn 20 points. For each half-court shot he misses, he will lose five points. After 20 half-court shots, Quentin has zero points. How many half-court shots did he make?
Answer by math-vortex(648)

Hi, there--

THE PROBLEM:
Quentin was challenged to a half-court shooting completion. For every half-court shot he makes, he will 
earn 20 points. For each half-court shot he misses, he will lose five points. After 20 half-court shots, 
Quentin has zero points. How many half-court shots did he make?

A SOLUTION:
Let x be the number of half-court shots that Quentin makes.

Quentin has 20 attempts all together, so the number he misses will be 20 minus the number he makes,
or 20-x.

He earns 20 points for each shot he makes, or 20x all together.

He loses 5 points for every shot he misses, or -5(20-x) all together.

After 20 attempts, the sum of his points is 0, so

20x - 5(20-x) = 0

Solve for x to find the number of half court shots he made. Distribute the 5 to get rid of the parentheses.

20x - 100 + 5x = 0

Combine like terms. (20x+5x is 25x.)
25x - 100 = 0

Add 100 to both sides to combine constant terms. (-100+100 is 0.)
25x = 100

Divide both sides by 25 to isolate the variable. (25x/25 is x.)
25x/25 = 100/25
x = 4

In the context of this problem, x = 4 means that Quentin made 4 half-court shots.


Let's make sure that answer makes sense. Quentin gets 20 points for every shot he makes. So he earned 
4*20=80 points. That left 16 misses because 20-4=16. He lost 5 points every time he missed. That's a 
loss of 80 points because 16*5= 80. He earned 80 points and lost 80 points. After 2 attempts, his score is 0.


Hope this helps! Feel free to email if you have any questions about the solution.

Good luck with your math,

Mrs. F
math.in.the.vortex@gmail.com